# On the Dynamic Regret of Following the Regularized Leader: Optimism with History Pruning

### v1

@@ Abstract
We revisit the Follow the Regularized Leader (FTRL) framework for Online Convex Optimization (OCO) over compact sets, focusing on achieving dynamic regret guarantees. Prior work has highlighted the framework’s limitations in dynamic environments due to its tendency to produce “lazy” iterates. However, building on insights showing FTRL’s ability to produce “agile” iterates, we show that it can indeed recover known dynamic regret bounds through optimistic composition of future costs and careful linearization of past costs, which can lead to pruning some of them. This new analysis of FTRL against dynamic comparators yields a principled way to interpolate between greedy and agile updates and offers several benefits, including refined control over regret terms, optimism without cyclic dependence, and the application of minimal recursive regularization akin to AdaFTRL. More broadly, we show that it is not the lazy projection style of FTRL that hinders (optimistic) dynamic regret, but the decoupling of the algorithm’s state (linearized history) from its iterates, allowing the state to grow arbitrarily. Instead, pruning synchronizes these two when necessary.
@@ 1.1 Background and Motivation
where $\tilde{\bm{g}}_{t}\in\partial\tilde{f}_{t}(\bm{x}_{t})$. Clearly, we can choose $\tilde{f}_{t}(\cdot)$ to be 00 or $f_{t-1}(\cdot)$, recovering the dependence on $G_{T}$ and $V_{T}$ [^1] More precisely, for differentiable functions, we recover $V^{\prime}_{T}=\sum_{t=1}^{T}\|\nabla{f}_{t}(\bm{x}_{t})-\nabla{f}_{t-1}(\bm{x}_{t-1})\|^{2}\leq V_{T}$., respectively. Algorithms whose regret depends on $E_{T}$ are called optimistic and are crucial for achieving best-of-both-worlds-style guarantees: constant regret in predictable environments and sub-linear in all cases, delivering adaptability without sacrificing robustness. Interestingly, the application of optimistic algorithms extends beyond enabling the use of untrusted predictions in the OCO problem; they have been shown to be key in the more general delayed OCO problem (Flaspohler et al., 2021), as well as the related OCO with memory problem (Mhaisen & Iosifidis, 2024). Given their significance, we focus on developing “optimistic” algorithms in this work. That is, algorithms that receive and use predictions of future costs and have regret bounds parametrized by $E_{T}$.
@@ 1.2 Methodology and Contributions
$\bm{x}_{t+1}=\operatorname*{argmin}_{\bm{x}}\langle{\bm{p}_{1:t}},{\bm{x}}\rangle+r_{1:t}(\bm{x})\\ +\tilde{f}_{t+1}(\bm{x})+I_{\mathcal{X}}(\bm{x}),$ (5)
@@
$\mathcal{N}_{\mathcal{X}}(\bm{x})\doteq\{\bm{g}\big{|}\langle{\bm{g}},{\bm{y}-\bm{x}}\rangle\leq 0,\forall\bm{y}\in\mathcal{X}\}.$
The simple yet key observation is that with the extra flexibility provided by $\bm{g}_{t}^{I}$, the state of the algorithm need not be the simple aggregation $\bm{g}_{1:t}$. Rather, some of the summands can be attenuated or pruned by carefully selecting $\bm{g}_{t}^{I}$ from the cone $\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$, This is possible since $\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$ contains all (scaled) negative subgradients of the expression in (5) when $\bm{x}_{t}$ lies on the boundary of $\mathcal{X}$ $\bm{x}_{t}\in\textbf{bd}(\mathcal{X})$ (from the optimality conditions for constrained problems, see e.g., (Beck, 2017, Thm. 3.67)). In words, when an iterate leaves the feasible set and is thus projected back, we can choose to prune the state that led to this situation and replace it with an alternative state $\bm{p}_{1:t}$, that induces the same iterate but is smaller in norm. This construction is crucial, as we later show that the norm of the state is the key bottleneck for FTRL when competing with time-varying comparators. Fig. 1 illustrates how different state constructions behave upon a switch in cost direction; note that the FTRL update, with a fixed $r(\cdot)$ can be expressed as the projection of $\nabla r^{\star}(-\bm{g}_{1:t})$, where $r^{\star}(\cdot)$ is the conjugate of $r(\cdot)$ (see def. in Appendix A.4).
Figure 1: Effect of dual state size on iterate agility. We consider two time steps $t<t^{\prime}$, where gradients switch direction: $\bm{g}_{\tau}=(-1,0)$ for $\tau\leq t$, and $\bm{g}_{\tau}=(1,0)$ for $\tau>t$. Top: Standard FTRL accumulates a large dual state $\bm{g}_{1:t}$, and the update via $\nabla r^{\star}(\bm{g}_{1:t})$ becomes insensitive to the change in direction; both $\bm{g}_{1:t}$ and $\bm{g}_{1:t^{\prime}}$ map to the same iterate. Bottom: A well-maintained dual state $\bm{p}_{1:t}$ remains bounded, and hence its mapping stays close to $\mathcal{X}$, enabling $\bm{x}_{t^{\prime}}$ to start aligning quickly with the new better direction.
@@ 2 Related Work
11:     Compute $\bm{x}_{t+1}^{\text{uc}}$ by solving (5) in $\mathbb{R}^{n}$.
@@ 3 OptFPRL
where $\bm{x}_{t}^{\text{uc}}$ is the unconstrained iterate obtained by solving (5) in $\mathbb{R}^{n}$ (i.e., without the indicator function).
To see why we can always set $\bm{g}_{t}^{I}$ as such, note that when $\bm{x}_{t}^{\text{uc}}\notin\mathcal{X}$, then $\bm{x}_{t}\in\textbf{bd}(\mathcal{X})$ (the projection of a point outside a compact convex set lies on the boundary). Hence, $\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$ contains vectors other than 00. In particular, it contains $-(\bm{p}_{1:t-1}+\bm{\tilde{g}}_{t}+\sigma_{1:t-1}\bm{x}_{t})$ and in fact all positive multiples of this vector. This follows directly from the optimality condition for the constrained iterate in (5), (see, e.g., (Beck, 2017, Thm. 3.67)), and the definition of the normal cone. For the second case, 00 is always a valid subgradient of the indicator function since, in this case, $\bm{x}_{t}^{\text{uc}}=\bm{x}_{t}\in\mathcal{X}$.
@@ 3.1 Dynamic Regret of OptFPRL
Settings 1. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a compact, convex set such that $\|\bm{x}\|\leq R$ for all $\bm{x}\in\mathcal{X}$. Let $\{f_{t}(\cdot),\tilde{f}_{t}(\cdot)\}_{t=1}^{T}$ be any sequence of $L$-Lipschitz convex functions. Define the path length $P_{T}$ as in (2), the cumulative prediction error $E_{T}$ as in (3), and the hybrid term $H_{T}\doteq\sum_{t=1}^{T}\epsilon_{t}\|\bm{u}_{t+1}-\bm{u}_{t}\|$.
@@ 3.1.2 Regularization with Prior $P_{T}$ Knowledge
In many cases, $P_{T}$ can be provided to the algorithm a priori as a measure of the comparator’s complexity (e.g., (Besbes et al., 2015; Si-Salem et al., 2024)). That is, we wish to compete against any sequence whose path length is at most the provided value of $P_{T}$. With this given target, we can adjust the regularization to account for the expected nonstationarity and obtain better bounds, as outlined next:
@@ 3.1.3 Regularization with Unknown $P_{T}$
Remarks. Note that this bound resembles that of Theorem 3.2, except for the additional term $A_{T}$, which arises due to the potential non-monotonicity of the estimated quantity $\sqrt{\nicefrac{{E_{t}}}{{P_{t}}}}$. If this sequence were monotonic and non-decreasing, $A_{T}$ would vanish since $[t]^{+}$ would be empty. This case allows us to recover the bound of Theorem 3.2 without requiring prior knowledge of $P_{T}$. In contrast, in the worst-case scenario, where the sequence alternates direction at every round, we obtain $A_{T}=\mathcal{O}(\sqrt{E_{T}}(P_{T}+1))$ (see Appendix C.3.1), leading to the looser bound in Theorem 3.1. A comparable correction term to $A_{T}$ also appears in the optimistic OMD framework (Scroccaro et al., 2023, Remark 2.19), and likewise depends on monotonicity, reaching $\sqrt{E_{T}}P_{T}$ in the worst case. The above-mentioned bound is, however, more interpretable.
@@ 4 Tools for FTRL’s Dynamic Regret Analysis
$\bm{x}_{t+1}\doteq\operatorname*{argmin}_{\bm{x}}h_{0:t}(\bm{x})+\tilde{f}_{t+1}(\bm{x})$
@@
[…]
@@
For any $\text{For any $t$, }\bm{x}_{t}^{\text{uc}}\in\mathcal{X}\implies\|\bm{p}_{1:t}\|\leq\sigma_{1:t-1}R+\epsilon_{t}.$ (36)
@@
From the first-order optimality condition for unconstrained problems (e.g., (Beck, 2017, Thm. 3.63)), we know that 00 is an element of the subdifferential of (37) at $\bm{x}_{t}$. Thus, $\exists\bm{\tilde{g}}_{t}\in\partial f_{t}(\bm{x}_{t})$ such that
@@
=$\stackrel{{\scriptstyle(a)}}{{=}}\|\bm{p}_{1:t-1}+\bm{g}_{t}\|\stackrel{{\scriptstyle(b)}}{{=}}\|-\sigma_{1:t-1}\bm{x}_{t}-\bm{\tilde{g}}_{t}+\bm{g}_{t}\|$ (41)
@@
For any $\text{For any $t$, }\bm{x}_{t}^{\text{uc}}\notin\mathcal{X}\implies\|\bm{p}_{1:t}\|\leq\sigma_{1:t-1}R+\epsilon_{t}.$ (44)
@@
$\bm{p}_{1:t}$ $=\bm{p}_{1:t-1}\ +\ \overbrace{\bm{g}_{t}\ \underbrace{-\bm{p}_{1:t-1}-\bm{\tilde{g}}_{t}-\sigma_{1:t-1}\bm{x}_{t}}}_{\text{\normalsize$\bm{g}_{t}^{I}$}}^{\text{\normalsize$\bm{p}_{t}$}}$ (45)
$=\bm{g}_{t}-\bm{\tilde{g}}_{t}-\sigma_{1:t-1}\bm{x}_{t}.$ (46)
[…]
$\leq\|\bm{g}_{t}-\bm{\tilde{g}}_{t}\|+\sigma_{1:t-1}\|\bm{x}_{t}\|$ (48)
$\leq\epsilon_{t}+\sigma_{1:t-1}R.$ (49)
@@ 4.1 Regret Bound Derivation
$\stackrel{{\scriptstyle(b)}}{{\leq}}4\sqrt{2}R\sqrt{E_{T}}+\frac{R}{8}\sqrt{E_{T}}$ (56)
@@
$\stackrel{{\scriptstyle(c)}}{{\leq}}5.8R\sqrt{E_{T}}+\frac{\sqrt{E_{T}}}{2}P_{T}+H_{T}$ (58)
@@
Proof sketch of other Theorems. The proofs of the other theorems follow similar main steps, and are detailed in the Appendix. When the $\sigma$ parameter is normalized by $\sqrt{P_{T}}$, it can be seen from (55) that the $P_{T}$ term will also be divided by the same, recovering the result of Theorem 3.2.
@@ A.2 Regularizer
[…]
@@ A.3 Pruning
Recall that this is a valid choice because when $\bm{x}_{t}^{\text{uc}}\notin\mathcal{X}$, then $\bm{x}_{t}\in\textbf{bd}(\mathcal{X})$. Hence, $\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$ contains elements other than 00, and in particular $-(\bm{p}_{1:t-1}+\bm{\tilde{g}}_{t}+\sigma_{1:t-1}\bm{x}_{t})$. This is a direct result of the optimality condition for the constrained iterate (59), see, e.g., (Beck, 2017, Thm. 3.67). For the second case, 00 is always a valid subgradient of the indicator function since in this case $\bm{x}_{t}^{\text{uc}}=\bm{x}_{t}\in\mathcal{X}$.

### v2

@@ Abstract
We revisit the Follow the Regularized Leader (FTRL) framework for Online Convex Optimization (OCO) over compact sets, focusing on achieving dynamic regret guarantees. Prior work has highlighted the framework’s limitations in dynamic environments due to its tendency to produce “lazy” iterates. However, building on insights showing FTRL’s ability to produce “agile” iterates, we show that it can indeed recover known dynamic regret bounds through optimistic composition of future costs and careful linearization of past costs, which can lead to pruning some of them. This new analysis of FTRL against dynamic comparators yields a principled way to interpolate between greedy and agile updates and offers several benefits, including refined control over regret terms, optimism without cyclic dependence, and the application of minimal recursive regularization akin to AdaFTRL. More broadly, we show that it is not the “lazy” projection style of FTRL that hinders (optimistic) dynamic regret, but the decoupling of the algorithm’s state (linearized history) from its iterates, allowing the state to grow arbitrarily. Instead, pruning synchronizes these two when necessary.
@@ 1.1 Background and Motivation
where $\tilde{\bm{g}}_{t}\in\partial\tilde{f}_{t}(\bm{x}_{t})$. Clearly, we can choose $\tilde{f}_{t}(\cdot)$ to be 0 or $f_{t-1}(\cdot)$, recovering the dependence on $G_{T}$ and $V_{T}$ [^1] More precisely, for differentiable functions, we recover $V^{\prime}_{T}=\sum_{t=1}^{T}\|\nabla{f}_{t}(\bm{x}_{t})-\nabla{f}_{t-1}(\bm{x}_{t-1})\|^{2}\leq V_{T}$., respectively. Algorithms whose regret depends on $E_{T}$ are called optimistic and are crucial for achieving best-of-both-worlds-style guarantees: constant regret in predictable environments and sub-linear in all cases, delivering adaptability without sacrificing robustness. Interestingly, the application of optimistic algorithms extends beyond enabling the use of untrusted predictions in the OCO problem; they have been shown to be key in the more general delayed OCO problem (Flaspohler et al., 2021), as well as the related OCO with memory problem (Mhaisen & Iosifidis, 2024). Given their significance, we focus on developing “optimistic” algorithms in this work. That is, algorithms that receive and use predictions of future costs and have regret bounds parametrized by $E_{T}$.
@@ 1.2 Methodology and Contributions
$\bm{x}_{t+1}=\operatorname*{argmin}_{\bm{x}}\langle{\bm{p}_{1:t}},{\bm{x}}\rangle+r_{1:t}(\bm{x})\\ +\tilde{f}_{t+1}(\bm{x})+I_{\mathcal{X}}(\bm{x}),\ignorespaces\ignorespaces$ (6)
@@
$\mathcal{N}_{\mathcal{X}}(\bm{x})\doteq\{\bm{g}\big|\langle{\bm{g}},{\bm{y}-\bm{x}}\rangle\leq 0,\forall\bm{y}\in\mathcal{X}\}.$ (9)
The simple yet key observation is that with the extra flexibility provided by $\bm{g}_{t}^{I}$, the state of the algorithm need not be the simple aggregation $\bm{g}_{1:t}$. Rather, some of the summands can be attenuated or pruned by carefully selecting $\bm{g}_{t}^{I}$ from the cone $\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$, This is possible since $\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$ contains all (scaled) negative subgradients of the expression in (LABEL:eq:update-rule-paper) when $\bm{x}_{t}$ lies on the boundary of $\mathcal{X}$ $\bm{x}_{t}\in\textbf{bd}(\mathcal{X})$ (from the optimality conditions for constrained problems, see e.g., (Beck, 2017, Thm. 3.67)). In words, when an iterate leaves the feasible set and is thus projected back, we can choose to prune the state that led to this situation and replace it with an alternative state $\bm{p}_{1:t}$, that induces the same iterate but is smaller in norm. This construction is crucial, as we later show that the norm of the state is the key bottleneck for FTRL when competing with time-varying comparators. Fig. 1 illustrates how different state constructions behave upon a switch in cost direction; note that the FTRL update, with a fixed $r(\cdot)$ can be expressed as the projection of $\nabla r^{\star}(-\bm{g}_{1:t})$, where $r^{\star}(\cdot)$ is the conjugate of $r(\cdot)$ (see def. in Appendix A.4).
Figure 1: Effect of dual state size on iterates agility. We consider two slots $t<t^{\prime}$, where gradients switch direction: $\bm{g}_{\tau}\!=\!(-1,0)$ for $\tau\leq t$, and $\bm{g}_{\tau}\!=\!(1,0)$ for $\tau\!>\!t$. Top: Standard FTRL accumulates a large state $\bm{g}_{1:t}$, and the update via $\nabla r^{\star}(\bm{g}_{1:t})$ becomes insensitive to the change in direction; both $\bm{g}_{1:t}$ and $\bm{g}_{1:t^{\prime}}$ map to the same iterate. Bottom: A well-maintained state $\bm{p}_{1:t}$ remains bounded, and hence its mapping stays close to $\mathcal{X}$, enabling $\bm{x}_{t^{\prime}}$ to start aligning quickly with the new better iterate direction (1,0)(-1,0).
@@ 2 Related Work
11:  Compute $\bm{x}_{t+1}^{\text{uc}}$ by solving (LABEL:eq:update-rule-paper) without $I_{\mathcal{X}}(\cdot)$.
@@ 3 OptFPRL
where $\bm{x}_{t}^{\text{uc}}$ is the unconstrained iterate obtained by solving (LABEL:eq:update-rule-paper) in $\mathbb{R}^{n}$ (i.e., without the indicator function).
To see why we can always set $\bm{g}_{t}^{I}$ as such, note that when $\bm{x}_{t}^{\text{uc}}\notin\mathcal{X}$, then $\bm{x}_{t}\in\textbf{bd}(\mathcal{X})$ (the projection of a point outside a compact convex set lies on the boundary). Hence, $\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$ contains vectors other than 0. In particular, it contains $(\bm{p}_{1:t-1}+\bm{\tilde{g}}_{t}+\sigma_{1:t-1}\bm{x}_{t})$ and in fact all positive multiples of this vector. This follows directly from the optimality condition for the constrained iterate in (LABEL:eq:update-rule-paper), (see, e.g., (Beck, 2017, Thm. 3.67)), and the definition of the normal cone. For the second case, 0 is always a valid subgradient of the indicator function since, in this case, $\bm{x}_{t}^{\text{uc}}=\bm{x}_{t}\in\mathcal{X}$.
@@ 3.1 Dynamic Regret of OptFPRL
Settings 1. Let $\mathcal{X}\subset\mathbb{R}^{d}$ be a compact, convex set such that $\|\bm{x}\|\leq R$ for all $\bm{x}\in\mathcal{X}$. Let $\{f_{t}(\cdot),\tilde{f}_{t}(\cdot)\}_{t=1}^{T}$ be any sequence of $L$-Lipschitz convex functions. Define the path length $P_{T}$ as in (3), the cumulative prediction error $E_{T}$ as in (4), and the hybrid term $H_{T}\doteq\sum_{t=1}^{T-1}\epsilon_{t}\|\bm{u}_{t+1}-\bm{u}_{t}\|$.
@@ 3.1.2 Regularization with Prior $P_{T}$ Knowledge
In many cases, $P_{T}$ can be provided to the algorithm a priori as a measure of the comparator’s complexity (e.g., (Si-Salem et al., 2024; Yang et al., 2016; Besbes et al., 2015)). That is, we wish to compete against any sequence whose path length is at most the provided value of $P_{T}$. With this given target, we can adjust the regularization to account for the expected nonstationarity and obtain better bounds, as outlined next:
@@ 3.1.3 Regularization with Unknown $P_{T}$
Remarks. Note that this bound resembles that of Theorem 3.2, except for the additional term $A_{T}$, which arises due to the potential non-monotonicity of the estimated quantity $\sqrt{\nicefrac{{E_{t}}}{{P_{t}}}}$. If this sequence were monotonic and non-decreasing, $A_{T}$ would vanish since $t]^{+}$ would be empty. This case allows us to recover the bound of Theorem 3.2 without requiring prior knowledge of $P_{T}$. In contrast, in the worst-case scenario, where the sequence alternates direction at every round, we obtain $A_{T}=\mathcal{O}(\sqrt{E_{T}}(P_{T}+1))$ (see Appendix C.3.1), leading to the looser bound in Theorem 3.1. A comparable correction term to $A_{T}$ also appears in the optimistic OMD framework (Scroccaro et al., 2023, Remark 2.19), and likewise depends on monotonicity, reaching $\sqrt{E_{T}}P_{T}$ in the worst case. The above-mentioned bound is, however, more interpretable.
@@ 4 Tools for FTRL’s Dynamic Regret Analysis
$\bm{x}_{t+1}\doteq\operatorname*{argmin}_{\bm{x}}h_{0:t}(\bm{x})+\tilde{f}_{t+1}(\bm{x})$ (32)
@@
$\textbf{(I)}\leq\min\left(\frac{1}{2}\|\bm{g}_{t}-\bm{\tilde{g}}_{t}\|_{t-1,*}^{2},2R\epsilon_{t}\right)$ (35)
$\textbf{(II)}\leq\left(\|\bm{p}_{1:t}\|+R\sigma_{1:t}\right)\|\bm{u}_{t+1}-\bm{u}_{t}\|$ (36)
@@
$\text{For any $t$, }\bm{x}_{t}^{\text{uc}}\in\mathcal{X}\implies\|\bm{p}_{1:t}\|\leq\sigma_{1:t-1}R+\epsilon_{t}.$ (39)
@@
From the first-order optimality condition for unconstrained problems (e.g., (Beck, 2017, Thm. 3.63)), we know that 0 is an element of the subdifferential of (40) at $\bm{x}_{t}$. Thus, $\exists\bm{\tilde{g}}_{t}\in\partial f_{t}(\bm{x}_{t})$ such that
@@
$\stackrel{{\scriptstyle(a)}}{{=}}\|\bm{p}_{1:t-1}+\bm{g}_{t}\|\stackrel{{\scriptstyle(b)}}{{=}}\|-\sigma_{1:t-1}\bm{x}_{t}-\bm{\tilde{g}}_{t}+\bm{g}_{t}\|$ (44)
@@
$\text{For any $t$, }\bm{x}_{t}^{\text{uc}}\notin\mathcal{X}\implies\|\bm{p}_{1:t}\|\leq\sigma_{1:t-1}R+\epsilon_{t}.$ (47)
@@
$\bm{p}_{1:t}$ $=\bm{p}_{1:t-1}\ +\ \overbrace{\bm{g}_{t}\ \underbrace{-\bm{p}_{1:t-1}-\bm{\tilde{g}}_{t}-\sigma_{1:t-1}\bm{x}_{t}}}_{\text{\normalsize$\bm{g}_{t}^{I}$}}^{\text{\normalsize$\bm{p}_{t}$}}$
$=\bm{g}_{t}-\bm{\tilde{g}}_{t}-\sigma_{1:t-1}\bm{x}_{t}.$
$\text{ Thus, }\|\bm{p}_{1:t}\|$ $=\|\bm{g}_{t}-\bm{\tilde{g}}_{t}-\sigma_{1:t-1}\bm{x}_{t}\|$
$\leq\|\bm{g}_{t}-\bm{\tilde{g}}_{t}\|+\sigma_{1:t-1}\|\bm{x}_{t}\|$
$\leq\epsilon_{t}+\sigma_{1:t-1}R.$
@@ 4.1 Regret Bound Derivation
$\stackrel{{\scriptstyle(b)}}{{\leq}}4\sqrt{2}R\sqrt{E_{T}}+\frac{R}{8}\sqrt{E_{T}}\$(54)
@@
$\stackrel{{\scriptstyle(c)}}{{\leq}}5.8R\sqrt{E_{T}}+\frac{\sqrt{E_{T}}}{2}P_{T}+H_{T}\$(56)
@@
Proof sketch of the other theorems. The proofs of the other theorems follow similar main steps, and are detailed in the Appendix. When the $\sigma$ parameter is normalized by $\sqrt{P_{T}}$, it can be seen from (53) that the $P_{T}$ term will also be divided by the same, recovering the result of Theorem 3.2.
@@ A.2 Regularizer
where $\sigma_{t}$ is the strong convexity parameter that will be set in every version of the algorithm. Note that the sum $r_{1:t}(\cdot)$ is $\sigma_{1:t}$-strongly convex with respect to $\|\cdot\|$, or equivalently $1$-strongly convex with respect to $\|\cdot\|_{t}\doteq\sqrt{\sigma_{1:t}}\|\cdot\|$.
@@ A.3 Pruning
Recall that this is a valid choice because when $\bm{x}_{t}^{\text{uc}}\notin\mathcal{X}$, then $\bm{x}_{t}\in\textbf{bd}(\mathcal{X})$. Hence, $\mathcal{N}_{\mathcal{X}}(\bm{x}_{t})$ contains elements other than 0, and in particular $(\bm{p}_{1:t-1}+\bm{\tilde{g}}_{t}+\sigma_{1:t-1}\bm{x}_{t})$. This is a direct result of the optimality condition for the constrained iterate (57), see, e.g., (Beck, 2017, Thm. 3.67). For the second case, 0 is always a valid subgradient of the indicator function since in this case $\bm{x}_{t}^{\text{uc}}=\bm{x}_{t}\in\mathcal{X}$.
Due to the FTRL/OMD connections, detailed in the related work, the update rule can be made equivalent to the “one-step optimistic MD” formulation in (Joulani et al., 2020), under the consistent pruning in (61) using the flexible $q_{t}(\cdot)$ terms in their Ada-MD update.